# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

### v1

@@ 3.2 Thermal dark matter relic via co-annihilation and
In Fig. 5, we present the points that yield the correct DM relic density in the $\Delta M$-$M_{\rm DM}$ plane. In the left plot, we show the points that give the correct relic density when co-scattering processes are excluded. The relic density calculation considers only annihilation, co-annihilation, and decay processes. The color coding represents different values of $\sin\theta$, ranging from 0.5 to $5\times 10^{-4}$, as indicated in the plot. As observed, for $\sin\theta>0.05$, the results are consistent with previous findings [8]. For $\sin\theta\leq 0.05$, we find new parameter space shown in magenta and cyan color points. For small $\Delta M$, decay and inverse-decay processes are comparable to co-annihilation processes. For a typical SD mixing, represented by magenta-colored points, the cross-section decreases as we move from left to right due to the increasing DM mass. This decrease can be compensated by reducing $\Delta M$. For larger $\Delta M$ [e.g., $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$], the co-annihilation cross-section experiences greater suppression, causing DM to decouple earlier and resulting in a larger abundance. However, with decay terms ($\Gamma_{D}$) included, the DM relic density can be brought into the correct range. Table 1 shows that for $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$, the DM is overabundant in the absence of decay. In the presence of decay and inverse-decay processes in the Boltzmann equation, DM remains in equilibrium for a longer duration, reducing the relic abundance to the desired value. The evolution of yields both considering decay and without decay has been shown in Fig. 12 in Appendix A. In the absence of decay terms, a small $\Delta M$ [e.g., $B({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\star})$] delays the DM decoupling epoch, resulting in a smaller yield ($Y_{1}(B)<Y_{1}(A)$). In this scenario, a smaller contribution from decay and inverse-decay processes is sufficient to achieve the correct relic density. This is evident in Table 1, where turning off the decay term results in a smaller increase in relic density compared to $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$. For even smaller $\Delta M$ [e.g., $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$], co-annihilation combined with decay-inverse decay fails to maintain equilibrium, leading to DM freezing out with a larger abundance. This behavior can be understood from $\Gamma_{D}\propto\sin^{2}\theta\Delta M^{5}$. For a fixed $\sin\theta$, as at point $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$, $\Delta M$ is too small to sustain equilibrium. Increasing $\sin\theta$ allows $\Gamma_{D}$ to adjust the relic density, as seen for $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$ with $\sin\theta=9.9\times 10^{-2}$. The magenta-colored points curve back leftward from $B({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\star})$. In this regime, decreasing DM mass and $\Delta M$ increases the cross-section, leading to a late decoupling of DM with a comparably smaller abundance. In this case, the contributions from decay terms are sufficient enough to provide the correct relic density. The turning point varies with $\sin\theta$: for smaller SD mixing, it occurs at smaller DM masses, while for larger SD mixing, it occurs at larger DM masses.
@@
[…]
@@
As discussed earlier, the co-scattering is important for small values of $\sin\theta$. In the right panel of Fig. 5, we present the correct relic density points, including the contribution from co-scattering processes. The parameter space that previously yielded the correct relic density when co-scattering was excluded gets modified due to additional processes. Conversely, the over-abundant points with $\sin\theta<10^{-4}$, which did not provide the correct relic density in the absence of co-scattering processes, now yield the correct relic density once these processes are included. For further clarification, we define a new parameter,
$\rho_{2}=\frac{\Omega_{2s}h^{2}}{\Omega_{2s}h^{2}({\rm no~{}coscattering})},$ (17)
@@
where $\Omega_{2s}h^{2}$ represents relic density, including co-scattering, while $\Omega_{2s}h^{2}$(no co-scattering) denotes relic density in the absence of the co-scattering contribution. From the definition of $\rho_{2}$ given in Eq. (17), $\rho_{2}=1$ indicates the relic density is independent of co-scattering effects. Consequently, any deviation from 1 will indicate the role of co-scattering in the final relic density. The larger is the deviation, the more is the effect from co-scattering. In fact, $\Omega_{2s}h^{2}$ includes all the processes involving annihilation, co-annihilation, co-scattering, and decay, while $\Omega_{2s}h^{2}$ (no co-scattering) involves only annihilation, co-annihilation, and decay. This implies the effective cross-section, giving rise to $\Omega_{2s}h^{2}$, is always larger than or equal to $\Omega_{2s}h^{2}$ (no co-scattering). Therefore, $\Omega_{2s}h^{2}\leq\Omega_{2s}h^{2}$ (no co-scattering), resulting in $\rho_{2}\leq 1$.
@@
Figure 6: Co-scattering parameter space in the plane of $\sin\theta$ vs $M_{\rm DM}$. The color code represents the parameter $\rho_{2}$.
@@
In Fig. 6, we present all the points that satisfy correct relic density in the plane of $\sin\theta$ vs $M_{\rm DM}$. The range of $\sin\theta$ is chosen in such a way that the DM will be in equilibrium. The color code depicts the $\rho_{2}$ values. If $\rho_{2}$=1, then co-scattering is sub-dominant, which is shown by the violet-colored points. As we decrease the value of $\sin\theta$, the $\rho_{2}$ deviates from 1, which implies the nontrivial role of co-scattering in large parameter space shown by cyan, yellow, and purple color points.
3.3 Freeze-in production of dark matter
@@
In Sections 3.1, and 3.2, we saw that the singlet is brought to equilibrium for the $\sin\theta\gtrsim\mathcal{O}(10^{-6})$. This implies that for $\sin\theta\gtrsim\mathcal{O}(10^{-6})$, the relic density of SDDM will be obtained via the freeze-out mechanism. If the SD mixing is further reduced, the DM relic density can no longer be determined through the conventional freeze-out mechanisms. However, it can be estimated through the freeze-in mechanism. In this regime, the heavier $\mathcal{Z}_{2}$-odd particles (belonging to sector 2) decay into DM (belonging to sector 1) in equilibrium, causing the DM number density to be produced via the freeze-in mechanism. The equation for the freeze-in mechanism, where particles from sector 2 decay into DM, is given by
$\frac{dY_{1}}{dx}=\frac{\Gamma_{tot}}{\mathcal{H}x}\frac{K_{1}(x)}{K_{2}(x)}Y^{eq}_{2},$ (18)
where $x=M_{\rm DM}/T$, and $\Gamma_{tot}$ is the total decay rate, which includes the two body decays like $\Gamma_{\chi_{0}\rightarrow\chi_{1}H},\Gamma_{\chi_{0}\rightarrow\chi_{1}Z},\Gamma_{\psi^{\pm}\rightarrow\chi_{1}W^{\pm}}$, and three body decay of the doublets as $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}l^{\pm}\nu_{l}}$ [^3] The decay mode $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}\pi^{\pm}}$ is very suppressed. This is because of $\Delta M\geq 1$ GeV.. The decay rates are given in Appendix B.
Figure 7: Parameter space for the correct relic density through freeze-in processes in the plane of $\sin\theta$ vs DM mass. The color code represents the mass splitting between the DM and next to the lightest $\mathcal{Z}_{2}$ odd particle.
In Fig. 7, we have shown the correct relic density points in the plane of $\sin\theta$ vs $M_{\rm DM}$ where the color code represents $\Delta M$. When $\Delta M\lesssim 80$ GeV, the freeze-in relic density is primarily determined by three-body decay modes. Conversely, as $\Delta M$ exceeds 80 GeV, 91 GeV, and 125 GeV, the two-body decay channels $\psi^{\pm}\rightarrow\chi_{1}W^{\pm}$, $\chi_{0}\rightarrow\chi_{1}Z$, and $\chi_{0}\rightarrow\chi_{1}H$ become accessible, respectively.
The three-body decay width is $\propto\sin^{2}\theta\Delta M^{5}$. When the $\Delta M$ increases, $\sin\theta$ has to be decreased to give the required decay width to produce the correct relic density. This trend is evident in Fig. 7. When the two body decay modes dominate in the larger mass-splitting regime, the decay width increases with an increase in $M_{\rm DM}$, $\Delta M$, and $\sin\theta$. For a fixed $\Delta M$, an increase in $M_{\rm DM}$ requires a reduction in SD mixing to achieve the correct relic density. Similarly, for a fixed $M_{\rm DM}$, an increase in $\Delta M$ necessitates adjusting the SD mixing to obtain the decay width that ensures the correct relic density. This behavior is illustrated in Fig. 7.
@@ 4 Direct detection of dark matter
$\sigma_{h}^{\rm SI}$ $=\frac{4}{\pi A^{2}}\mu^{2}_{r}\frac{\Delta M^{2}\sin^{4}2\theta}{2v^{2}M^{4}_{h}}\Big{[}\frac{m_{p}}{v}\Big{(}f^{p}_{Tu}+f^{p}_{Td}+f^{p}_{Ts}+\frac{2}{9}f^{p}_{TG}+\frac{m_{n}}{v}\Big{(}f^{n}_{Tu}+f^{n}_{Td}+f^{n}_{Ts}+\frac{2}{9}f^{n}_{TG}\Big{)}\Big{]}^{2}.$ (19)
Different coupling strengths between DM and light quarks are given by [42, 43] as $f^{p}_{Tu}=0.020\pm 0.004,~{}f^{p}_{Td}=0.026\pm 0.005,~{}f^{p}_{Ts}=0.014\pm 0.062,~{}f^{n}_{Tu}=0.020\pm 0.004,~{}f^{n}_{Td}=0.036\pm 0.005,~{}f^{n}_{Ts}=0.118\pm 0.062$. The coupling of DM with the gluons in target nuclei are parameterized by [44] $f^{p,n}_{TG}=1-\sum_{q=u,d,s}f^{p,n}_{Tq}$. Similarly, the SI DM-nucleon scattering cross-section through Z mediation is given by [45, 46]:
$\sigma^{\rm SI}_{Z}=\frac{G^{2}_{F}\sin^{4}\theta}{\pi A^{2}}\mu_{r}^{2}\Big{|}\left[Zf_{p}+(A-Z)f_{n}\right]^{2}\Big{|}^{2},$ (20)
@@
A larger mixing angle leads to an increased direct detection cross-section, which is excluded by the direct detection experiment like LZ [6]. We calculated the SI direct detection cross-section ($\sigma^{\rm SI}_{\rm DM-N}=\sigma_{h}^{\rm SI}+\sigma^{\rm SI}_{Z}$) for the correct relic points and shown w.r.t $M_{\rm DM}$ in Fig. 9. The direct detection excludes the SD mixing angle for $\sin\theta>0.07$. We have also shown the LZ exclusion region in the parameter space of $\sin\theta$ vs $M_{\rm DM}$ in Fig. 11. The DARWIN [47] sensitivity is shown with the blue dashed line. When we include the LEP bound, the exclusion goes to $\sin\theta>0.04$.

### v2

@@ 3.1.2 Thermal dark matter relic via the conversion-driven processes in the small $\sin\theta$ regime
To explain our findings, we identify five sample points: $A,B,C,D$ and $E$ as shown in the leftleft panel of Fig. 6. For a typical SD mixing, represented by magenta-colored points, the cross-section decreases as we move from left to right due to the increasing DM mass. This decrement can be compensated by reducing $\Delta M$. In absence of decay and inverse decay terms in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), for larger $\Delta M$ [e.g., point $A$], the co-annihilation cross-section experiences greater suppression, causing DM to decouple earlier and resulting in a larger abundance. However, with decay terms ($\Gamma_{\Psi\rightarrow\chi,\rm SM}$) included, the DM relic density can be brought to the correct range. Table 1 shows that for point $A$, the DM is overabundant in absence of decay term [^6] In Table 1, last column corresponds to the relic without considering co-scattering and decay, which we denote as $\Omega_{2s}h^{2}$(no co-scattering) without decay. This can be achieved using the function darkOmegaN with the flag ExcludedForNDM="2010 DMdecay". On the other hand $\Omega_{2s}h^{2}$(no co-scattering) is obtained using the function darkOmegaN with the flag ExcludedForNDM="2010".. In presence of the decay and inverse-decay processes in the Boltzmann equations (LABEL:eq:Y1) and (LABEL:eq:Y2), sector 1 to sector 2 conversion can happen efficiently for a longer duration, thereby reducing the relic abundance to the desired value. The evolution of yields ($Y_{1}$ and $Y_{2}$) both considering decay and without decay has been shown in Fig. 15 in Appendix A. In absence of the decay terms, a relatively small $\Delta M$ [e.g., point $B$] delays the DM decoupling epoch, resulting in a smaller yield in comparison to point AA ($Y_{1}(B)<Y_{1}(A)$). In this scenario, a smaller contribution from decay and inverse-decay processes is sufficient to achieve the correct relic density. This is evident from Table 1, where turning off the decay term results in a smaller increase in relic density compared to point $A$. For further smaller $\Delta M$ [e.g., point $C$ where $\sin\theta=9.5\times 10^{-3}$], co-annihilation combined with decay-inverse decay fail to maintain equilibrium, leading to DM freezing out with a larger abundance in comparison to point $B$. This behavior can be understood from the decay term: $\Gamma_{\Psi\rightarrow\chi,\rm SM}\propto\sin^{2}\theta\Delta M^{5}$. Since $\sin\theta$ is same for point $B$ and $C$, the reduced $\Delta M$ at the point $C$ gives a smaller $\Gamma$, thereby resulting a larger relic. By increasing $\sin\theta$ from $9.5\times 10^{-3}$ to $9.9\times 10^{-2}$ we recover the correct relic density as shown in Table 1.
For smaller mixing angles, typically $\sin\theta<0.05$, we have new sets of point that satisfy correct relic density. In particular, the set of points represented by magenta and cyan color curve back leftward from point $B$ and $E$, respectively. In this regime, decreasing DM mass and $\Delta M$, increase the cross-section, leading to a late decoupling of DM with a comparatively smaller abundance yet larger than the required relic. In this case, the small contribution from decay terms are required to give rise the correct relic density. The position of the turning points $B$ and $E$ varies with $\sin\theta$. For smaller SD mixing, it occurs at smaller DM masses, while for larger SD mixing, it occurs at larger DM masses. This can be understood as follows. The cross section usually decreases with increase in mass, leading to a larger chemically decoupled relic. Therefore, for a larger DM mass, we need a larger $\sin\theta$ to bring down the relic to the correct ballpark and vice-versa.
@@
AA 100 8.5 $6\times 10^{-3}$ 0.1128 $4.795\times 10^{2}$
$B$ 650 3 $9.5\times 10^{-3}$ 0.1269 $2.503\times 10^{1}$
$C$ 700 1.5 $9.5\times 10^{-3}$ 0.3445 $2.114\times 10^{1}$
$D$ 700 1.5 $9.9\times 10^{-2}$ 0.1234 0.1434
$E$ 520 6.2 $10^{-3}$ 0.1247 $2.23\times 10^{3}$
@@
As discussed earlier, for larger SD mixing ($\sin\theta\gtrsim 0.05$) the sector 1 (singlet) and sector 2 (doublet) decouple almost at the same epoch. In this case, the co-scattering effects does not play vital role in deciding the relic density of DM, even though the rate of co-scattering processes are large. This implies our results are consistent with previous findings [6]. However, for smaller SD mixing ($\sin\theta<0.05$), the co-scattering effects are important. In the right panel of Fig. 6, we present the correct relic density points by solving the Boltzmann equations (LABEL:eq:Y1) and (LABEL:eq:Y2), without excluding the contribution from co-scattering processes. The parameter space that previously yielded the correct relic density when co-scattering was excluded get under abundant in presence of the co-scattering processes. Conversely, the over-abundant points with $\sin\theta<10^{-4}$, which did not provide the correct relic density in the absence of co-scattering processes, now yield the correct relic density once these processes are included. This can be understood by examining Eq. (3.8), which includes both co-scattering as well as decay and inverse decay processes. In the small $\sin\theta$ limit, the co-scattering rates are much larger than the decay and inverse decay rates (see right panel of Fig. 3), even though they have same $\sin\theta$ dependency. Therefore, for a given $\sin\theta$, which yields the correct relic density in the absence of co-scattering, become under abundant in presence of co-scattering processes as the latter processes will convert the sector 1 to sector 2 particles followed by annihilation to SM bath in a faster rate. We consolidate our claim by defining a new parameter, $\rho_{2}$, as
$\rho_{2}=\frac{\Omega_{2s}h^{2}}{\Omega_{2s}h^{2}({\rm no\leavevmode\nobreak\ coscattering})},$ (3.12)
@@
where $\Omega_{2s}h^{2}$ represents relic density obtained by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), while $\Omega_{2s}h^{2}$(no co-scattering) denotes the solution of Eqs.(LABEL:eq:Y1) and (LABEL:eq:Y2) in the absence of the co-scattering contribution. Eq. (3.12) implies that, $\rho_{2}=1$ indicates the relic density is independent of co-scattering effects. Consequently, any deviation from 1 will indicate the role of co-scattering in the final relic density. The larger is the deviation, the more is the effect from co-scattering. In fact, $\Omega_{2s}h^{2}$ includes all the processes involving annihilation, co-annihilation, co-scattering, and decay, while $\Omega_{2s}h^{2}$ (no co-scattering) involves only annihilation, co-annihilation, and decay. This implies the effective cross-section, giving rise to $\Omega_{2s}h^{2}$, is always larger than or equal to $\Omega_{2s}h^{2}$ (no co-scattering). Therefore, $\Omega_{2s}h^{2}\leq\Omega_{2s}h^{2}$ (no co-scattering), resulting in $\rho_{2}\leq 1$.
@@
Figure 7: The co-scattering parameter space in the plane of $\sin\theta$ vs $M_{\rm DM}$. The color code represents the parameter $\rho_{2}$.
@@
In Fig. 7, we present all the points that satisfy correct relic density in the plane of $\sin\theta$ vs $M_{\rm DM}$. The range of $\sin\theta$ is chosen in such a way that the DM will be in equilibrium. The color code depicts the $\rho_{2}$ values. If $\rho_{2}$=1, then the effect of co-scattering is negligible, which is shown by the violet-colored points. As we decrease the value of $\sin\theta$, the $\rho_{2}$ deviates from 1, which implies the nontrivial role of co-scattering in large parameter space shown by cyan, yellow, and purple color points. Below this region, the DM is produced thermally. However the conversion-driven processes (both decay, inverse decay and co-scattering) are not efficient enough due to further smaller $\sin\theta$. As a result, in this region, the DM remains over abundant. Beyond certain threshold limit of SD mixing, the DM (dominantly singlet component) never comes to thermal equilibrium. In this region of parameter space, the DM can be produced via freeze-in processes as we discussed below.
3.2 Dark matter production via freeze-in mechanism
@@
In Sections 3.1.1, and 3.1.2, we saw that the singlet is brought to equilibrium for the $\sin\theta\gtrsim\mathcal{O}(10^{-7})$ depending on the DM mass. This implies that for $\sin\theta\gtrsim\mathcal{O}(10^{-7})$, the relic density of SDDM will be obtained via the freeze-out mechanism. If the SD mixing is further reduced, the DM relic density can no longer be determined through the conventional freeze-out mechanisms. However, it can be produced through the freeze-in mechanism. In this regime, we take into account all the $2\rightarrow 2$ processes along with decay and inverse decay processes and solve Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) with initial conditions, $Y_{2}=Y_{2}^{\rm eq}$ and $Y_{1}=0$. To estimate the final DM relic density, we used $\Omega_{\rm DM}h^{2}=0.12\left(\frac{(Y_{1}+Y_{2})_{T\rightarrow 0}}{4.26\times 10^{-10}}\right)\left(\frac{M_{\rm DM}}{{1\leavevmode\nobreak\ \rm GeV}}\right)$.
Figure 8: Variation of relic density with DM mass for a fixed $\Delta{M}=1$ GeV. We take 5 values of $\sin\theta$ as shown by different colors. The black dashed line represents the required relic density.
Figure 9: Contours of the correct relic density obtained through freeze-in processes are shown by the colored solid lines in the plane of $\sin\theta$ vs DM mass. The corresponding BBN constraints are shown by colored dashed lines.
In Fig. 8, we present discrete contours of $\sin\theta$ in the $\Omega_{\rm DM}h^{2}-M_{\rm DM}$ plane, assuming a fixed SD mass splitting $\Delta M=1$ GeV. We observe that increasing $\sin\theta$ enhances the rates of both $2\rightarrow 2$ scattering and decay-induced production processes, leading to a higher dark matter relic abundance. Additionally, for each $\sin\theta$ the relic density increases with increasing $M_{\rm DM}$. This behavior arises because, in the freeze-in production scenario, the dark matter yield saturates around $T_{\rm FI}\sim M_{\rm DM}$. A higher freeze-in temperature $(T_{\rm FI})$ corresponds to a larger production rate, thereby resulting in a larger relic abundance for larger $M_{\rm DM}$.
In Fig. 9, we have shown the correct relic density contours with solid lines for five different mass splittings, $\Delta{M}=1,10,100,500,1000$ GeV, in the plane of $\sin\theta$ verses $M_{\rm DM}$. We also show the corresponding BBN constraints [^7] We take a conservative time scale of $\tau_{BBN}=1$ s. In the region below the dashed lines, $\tau_{BBN}<\tau_{\psi}$, and hence ruled out. with dashed lines, arising from the doublet decay to the singlet. Since these contours (solid lines) correspond to correct relic density, with increase in mass, the number density is expected to decrease [^8] As $\Omega_{\rm DM}h^{2}\propto M_{\rm DM}n_{\rm DM}$, with increase in mass of DM, the number density of DM should decrease to provide relic density in the correct ballpark. It is possible if we decrease the $\sin\theta$ which leads to a reduction in the DM number density., which requires a smaller $\sin\theta$. This behavior can be easily read from the solid contours shown in Fig. 9. When $\Delta M\lesssim 80$ GeV, the freeze-in relic density is primarily determined by three-body decay modes of the doublet along with other $2\rightarrow 2$ processes like $22\rightarrow 11,20\rightarrow 10$, etc. Conversely, as $\Delta M$ exceeds 80 GeV, 91 GeV, and 125 GeV, the two-body decay channels $\psi^{\pm}\rightarrow\chi_{1}W^{\pm}$, $\chi_{0}\rightarrow\chi_{1}Z$, and $\chi_{0}\rightarrow\chi_{1}H$ become accessible along with the above $2\rightarrow 2$ processes. The red solid line represents $\Delta{M}=10$ GeV. The corresponding BBN constraint excludes this $\Delta{M}$ for all range of DM masses. On the other hand, BBN excludes $M_{DM}>60$ GeV for $\Delta{M}=100$ GeV which is shown with the magenta colored line. The green colored line represents $\Delta{M}=500$ GeV. We see that the correct relic can be achieved only up to $M_{\rm DM}\sim 420$ GeV. Beyond this range of DM mass, the doublet mass becomes >920>920 GeV. A pure doublet fermion DM achieves the correct relic near $M_{\psi}\sim 1000$ GeV. In this case, the doublet decay alone produces nearly the required relic while the presence of other $2\rightarrow 2$ processes enhances the production leading to an over abundant relic. Therefore, the region right to the contours is excluded as the DM becomes overabundant in this region. For the same reason, in case $\Delta{M}=1000$ GeV, the maximum DM mass allowed by the relic density constraint is $\sim 200$ GeV. Another observation from Fig. 9, is that for DM mass <1<1 TeV with $\Delta{M}>1$ GeV, the correct relic of DM is obtained for $\sin\theta$ varying in the range $10^{-13}-10^{-11}$} and are not excluded by the BBN constraint. For $10^{-7}>\sin\theta>10^{-11}$, the DM is over produced through freeze-in processes.
3.3 Dark matter production via SuperWIMP mechanism
For $\sin\theta<\mathcal{O}(10^{-14})$, the DM relic can be achieved using the SuperWIMP mechanism [47, 48, 15]. In this range of $\sin\theta$, $2\rightarrow 2$ productions are negligible and DM gets produced from the out-of-equilibrium decay of the doublet. The DM relic can be expressed in terms of the doublet relic as
$\Omega_{\rm DM}h^{2}=\left(\frac{M_{\rm DM}}{M_{\Psi}}\right)\Omega_{\Psi}h^{2},$ (3.13)
Figure 10: Left: Relic abundance of $\Psi$ as function of $M_{\Psi}$. Right: Correct relic parameter space obtained via the SuperWIMP mechanism is shown in the plane of $\Delta{M}-M_{\rm DM}$ with the blue line. The colored dashed lines represents the BBN exclusion limits for different $\sin\theta$ as mentioned in the plot.
where $\Omega_{\Psi}h^{2}$ is the relic of the doublet.
In the left panel of Fig. 10, we show the relic of the doublet as a function of $M_{\Psi}$. In the right panel of Fig. 10, correct DM relic parameter space is shown in the plane of $\Delta{M}-M_{\rm DM}$ where the $\Delta{M}$ is varied in the range $10^{3}-10^{5}$} GeV. Utilizing Eq. (3.13), we calculated the values of $M_{\rm DM}$ that satisfies the correct relic density of the DM. We have also shown the BBN constraints on the decay of the doublet in the plane of $\Delta{M}-M_{\rm DM}$ for 3 different choices of $\sin\theta$ as $10^{-14}$ (red dashed), $10^{-15}$ (dark magenta dashed) and $10^{-16}$ (green dashed). The region below each dashed lines are excluded from the BBN for the corresponding $\sin\theta$.
@@ 4 Direct detection of dark matter
$\sigma_{h}^{\rm SI}$ $=\frac{4}{\pi A^{2}}\mu^{2}_{r}\frac{\Delta M^{2}\sin^{4}2\theta}{2v^{2}M^{4}_{h}}\Big{[}\frac{m_{p}}{v}\Big{(}f^{p}_{Tu}+f^{p}_{Td}+f^{p}_{Ts}+\frac{2}{9}f^{p}_{TG}+\frac{m_{n}}{v}\Big{(}f^{n}_{Tu}+f^{n}_{Td}+f^{n}_{Ts}+\frac{2}{9}f^{n}_{TG}\Big{)}\Big{]}^{2}.$ (4.1)
Different coupling strengths between DM and light quarks are given by [55, 56] as $f^{p}_{Tu}=0.020\pm 0.004,\leavevmode\nobreak\ f^{p}_{Td}=0.026\pm 0.005,\leavevmode\nobreak\ f^{p}_{Ts}=0.014\pm 0.062,\leavevmode\nobreak\ f^{n}_{Tu}=0.020\pm 0.004,\leavevmode\nobreak\ f^{n}_{Td}=0.036\pm 0.005,\leavevmode\nobreak\ f^{n}_{Ts}=0.118\pm 0.062$. The coupling of DM with the gluons in target nuclei are parameterized by [57] $f^{p,n}_{TG}=1-\sum_{q=u,d,s}f^{p,n}_{Tq}$. Similarly, the SI DM-nucleon scattering cross-section through Z mediation is given by [58, 59]:
$\sigma^{\rm SI}_{Z}=\frac{G^{2}_{F}\sin^{4}\theta}{\pi A^{2}}\mu_{r}^{2}\Big{|}\left[Zf_{p}+(A-Z)f_{n}\right]^{2}\Big{|}^{2},$ (4.2)
@@
A larger mixing angle leads to an increased direct detection cross-section, which is excluded by the direct detection experiment like LZ [7]. We calculated the SI direct detection cross-section ($\sigma^{\rm SI}_{\rm DM-N}=\sigma_{h}^{\rm SI}+\sigma^{\rm SI}_{Z}$) for the correct relic points and shown w.r.t $M_{\rm DM}$ in Fig. 12. We have also shown the LZ exclusion region in the parameter space of $\sin\theta$ vs $M_{\rm DM}$ in the same plot. It excludes the SD mixing angle for $\sin\theta>0.07$. The DARWIN [60] sensitivity is shown with the blue dashed line. When we include the LEP bound, the exclusion goes to $\sin\theta>0.04$.